OpenStudy (anonymous):

Solve. 8x – 11 + 6x = –151

OpenStudy (anonymous):

Add the like terms. 8x + 6x = 14x 14x - 11 = -151 Add 11 to each side. -151 + 11 = -140 14x = -140 -140 divided by 14 = -10 x = -10